# A Survey on Neural Topic Models: Methods, Applications, and Challenges

### v1

@@ 3.3 NTMs with Metadata
While common NTMs learn topics in an unsupervised manner (only using documents), NTMs can also leverage the metadata of documents to guide topic modeling, similar to supervised LDA (Mcauliffe and Blei, 2007). In detail, Card et al. (2018) introduce a NTM that can incorporate various metadata of documents. It encodes a document together with its labels (e.g., sentiment) and covariates (e.g., publication year), and generates the document conditioned on the covariates. Korshunova et al. (2019) model the generation of documents and labels together in a discriminative way; then train their model with mean-field variational inference. They can also incorporate a variety of data modalities like images. Wang and Yang (2020) jointly model topics and train a RNN classifier to predict document labels. They are connected by an attention mechanism. Wang et al. (2021a) incorporate document networks in a NTM and jointly reconstruct documents and networks.
@@ 3.6 NTMs with Pre-trained Language Models
Researchers frequently combine NTMs with pre-trained language models. Pre-trained language models based on Transformers (Vaswani et al., 2017) have been prevalent in NLP fields, which are pre-trained on large-scale corpora to capture contextual linguistic features. Multiple studies leverage contextual features from these pre-trained models to provide richer information than conventional BoW. For instance, Bianchi et al. (2021a) input the concatenation of BoW and the contextual document embeddings from Sentence-BERT (Reimers and Gurevych, 2019), and then reconstruct BoW as previous work. Hoyle et al. (2020) propose to distill knowledge from BERT (Devlin et al., 2018) to NTMs. In detail, they produce pseudo BoW from the predictive word probability of BERT. Then their NTM reconstructs both the real and pseudo BoW. Bianchi et al. (2021b); Mueller and Dredze (2021) employ multilingual BERT to infer cross-lingual doc-topic distributions for zero-shot learning but they cannot discover aligned cross-lingual topics.
@@ 3.7 NTMs with Contrastive Learning
As a self-supervised learning fashion, contrastive learning has been employed to facilitate NTMs. The idea of contrastive learning is to measure the similarity relations among sample pairs in a representation space (Van den Oord et al., 2018). Nguyen and Luu (2021) propose the contrastive learning on doc-topic distributions where they build positive and negative pairs by sampling salient words of documents. Differently, Wu et al. (2022) directly sample positive and negative pairs based on the topic semantics of documents to capture relations among samples. Specifically, they quantize doc-topic distributions following Wu et al. (2020b) and then sample documents with the same quantization indices as positive pairs and different indices as negative pairs. Their method can also capture the similarity relations among additional augmented data. Zhou et al. (2023) improve topic disentanglement with contrastive learning on word and topic embeddings. Han et al. (2023) cluster documents, compute term weights, and make NTMs reconstruct salient words. They also use contrastive learning to refine doc-topic distributions where positive samples come from pre-trained language models.
@@ 3.9 Other NTMs
Before the invention of VAE-based NTMs, researchers have different attempts to model latent topics with neural networks. Some studies focus on NTMs in the autoregressive framework. Larochelle and Lauly (2012) propose an autoregressive NTM, called DocNADE. Inspired by Replicated Softmax (Hinton and Salakhutdinov, 2009), DocNADE predicts the probability of a word in a document conditioned on its hidden state which is conditioned on previous words. Then it interprets topics with a hidden state and infers doc-topic distributions with the hidden states of the document. Gupta et al. (2019b) extend DocNADE by modeling the bi-directional dependencies between words. Gupta et al. (2019a) then use a LSTM to enable DocNADE to incorporate external knowledge.
Cao et al. (2015) also propose an early NTM before VAE-based NTMs. Their approach predicts how an n-gram correlates with documents. It computes the representation of an n-gram by transforming the accumulation of the word embeddings from Word2Vec (Mikolov et al., 2013) and projects documents into representations with a look-up matrix table. In this way, they model topic-word distributions as the n-gram representations and model doc-topic distributions as the projected document representations. For training, it uses the document of the n-gram as a positive and randomly samples documents that do not contain this n-gram as negatives.
@@ 3.10 Topic Discovery by Clustering
We discuss a special type of approach that discovers latent topics by clustering instead of modeling the generation process of documents. They typically leverage traditional word embeddings such as Word2Vec (Mikolov et al., 2013) or contextual embeddings from pre-trained language models. We must clarify that these methods differ from the aforementioned ordinary topic models. This is because they can only produce topics but cannot infer doc-topic distributions as required. Accordingly, their one advantage is their enhanced computational efficiency. In detail, Thompson and Mimno (2020) straightforwardly cluster token-level word embeddings from pre-trained models like BERT and GPT-2 and produce topics from the words assigned to clusters. Similarly, Sia et al. (2020); Angelov (2020); Zhang et al. (2022c) cluster word or document embeddings and interpret hidden topics by sampling words from clusters via term weights like TF-IDF. Following this line of work, Grootendorst (2022) propose BERTopic by clustering document representations through HDBSCAN. Note that BERTopic estimates the doc-topic distribution based on the term weights within a given document.
@@ 4.1 Hierarchical NTMs
Similar to conventional topic models (Griffiths et al., 2003; Teh et al., 2004; Blei et al., 2010), NTMs can discover hierarchical topics to reveal topic structures from general to specific. Topics at each level in a hierarchy cover different semantic granularity: child topics tend to be more specific to their parent topics. As shown in Figure 4, a topic about “sports” can derive more specific child topics, like “soccer”, “basketball”, and “tennis”; a topic about “computer” also has specific child topics like “linux”, “programming”, and “windows”. In addition, hierarchical topic modeling can relieve the challenge of determining the number of topics to some extent (Blei et al., 2010).
To discover hierarchical topics, some NTMs follow the previous non-parametric setting which allows topic hierarchies to grow dynamically. Isonuma et al. (2020) propose a tree-structured neural topic model with two doubly-recurrent neural networks over the ancestors and siblings respectively (Alvarez-Melis and Jaakkola, 2017). Note that the tree structure is unbounded, i.e., it can be dynamically updated in a heuristic way during training. Pham and Le (2021) follow this spirit and jointly handle hierarchical topics and document visualization. Chen et al. (2021b) leverage a stick-breaking process as prior for non-parametric modeling.
Later, the parametric fashion has attracted more attention, which presets the width and height of a topic hierarchy before learning. This leads to the main issue that topic hierarchies cannot grow dynamically. Chen et al. (2021a) propose manifold regularization on topic hierarchy learning. Duan et al. (2021) propose a Sawtooth Connection to model topic dependencies across hierarchical levels based on the model structure of ETM (Dieng et al., 2020). As aforementioned, Xu et al. (2022) use different layers in the hyperbolic embedding space to interpret hierarchical topics. Li et al. (2022) employ skip-connections for decoding to alleviate the posterior collapsing issue and propose a policy gradient method for training. Recently, Duan et al. (2023) propose to generate different documents for different levels. They craft documents with more related words through word similarity matrices for higher levels, and then progressively generate these documents at each level. Chen et al. (2023) utilize a Gaussian mixture prior and nonlinear structural equations to model topic dependencies between hierarchical levels. Wu et al. (2024) propose to produce sparse and balanced dependencies by modeling them as the transport plan solutions of specially defined optimal transport problems between hierarchical topics. They then introduce a context-aware disentangled decoder to decode documents with topics at each level, which better distributes different semantic granularity to different levels.
@@ 4.2 Short Text NTMs
Researchers apply NTMs to discover topics from short texts. Short texts, prevalent on the Internet in various forms such as tweets, comments, and news headlines, serve as a common medium for individuals to express ideas, comments, and opinions. However, normal topic models often struggle to effectively handle short texts. The principal reason is that topic models depend on the word co-occurrence information to infer latent topics, but such information is extremely sparse in short texts due to their limited context. This challenge, referred to as data sparsity (Yan et al., 2013; Wu and Li, 2019), hinders topic models from discovering high-quality topics and thus has attracted considerable attention in the research community.
Several studies are proposed to overcome this data sparsity challenge. Lin et al. (2020) use the Archimedean copulas to regularize the discreteness of topic distributions of short texts. Wu et al. (2020b) propose to quantize doc-topic distributions of short texts to quantization vectors following the idea of Van den Oord and Vinyals (2017). By carefully initializing the quantization vectors, they can produce sharper doc-topic distributions that better fit short texts with limited context. They also propose a negative sampling decoder to avoid repetitive topics besides the negative log-likelihood. To address the data sparsity issue, Wang et al. (2021b) use word co-occurrence and semantic correlation graphs to enrich the learning signals of short texts. Zhao et al. (2021d) propose to incorporate entity vector representations into a NTM for short texts. They learn entity vector representations from manually edited knowledge graphs. Based on Wu et al. (2020b), Wu et al. (2022) further propose a contrastive learning method according to the topic semantics of short texts, which better captures the similarity relations among them. This refines the representations of short texts and thus their doc-topic distributions. They can also adapt to using data augmentation to further mitigate the data sparsity problem.
@@ 4.3 Cross-lingual NTMs
Cross-lingual NTMs are proposed following cross-lingual topic models (Mimno et al., 2009). Cross-lingual topic models aim to discover aligned topics in different languages. As exemplified in Figure 5, English and Chinese Topic#3 both refer to “music”, and English and Chinese Topic#5 refer to “celebrity”. In addition if two documents in different languages contain similar latent topics, their inferred doc-topic distributions should be similar. For instance, the doc-topic distributions of the parallel English and Chinese documents in Figure 5 are similar. These aligned cross-lingual topics can reveal commonalities and differences across languages and cultures, which enables cross-lingual text analysis without supervision.
Wu et al. (2020a) propose the first neural cross-lingual topic model. It transforms the topic-word distribution to the vocabulary space of another language. Thus the topic-word distributions of one language can incorporate the semantics of another language, which aligns cross-lingual topics. They show that their model outperforms traditional multilingual topic models (Shi et al., 2016; Yuan et al., 2018). Later, Wu et al. (2023a) propose to align cross-lingual topics from the perspective of mutual information. This can properly align cross-lingual topics and prevent degenerate topic representations. To address the low-coverage dictionary issue, they also propose a cross-lingual vocabulary linking method that finds more linked words for topic alignment beyond the given dictionary. Bianchi et al. (2021b); Mueller and Dredze (2021) directly learn cross-lingual doc-topic distributions with multilingual BERT. But we emphasize that they cannot discover aligned cross-lingual topics as required.
@@ 4.4 Dynamic NTMs
Dynamic NTMs are explored following dynamic topic models (Blei and Lafferty, 2006b; Wang et al., 2008). Previous static topic models implicitly assume that documents are exchangeable. However, this assumption is inappropriate since documents are produced sequentially, such as scholarly journals, emails, and news articles. As such, dynamic topic models are proposed. While topics in previous methods are all static, dynamic topic models allow topics to shift over time to capture the topic evolution in sequential documents.
@@
Dieng et al. (2019) first propose a neural dynamic topic model, the Dynamic Embedding Topic Model. It uses word and topic embeddings to interpret latent topics following Dieng et al. (2020) and chains topic embeddings at slice $t$ with topic embeddings at slice $t-1$ by Markov chains. Besides, it uses a LSTM to learn temporal priors of doc-topic distributions. Rahimi et al. (2023) discover topic evolution by clustering documents but cannot infer doc-topic distributions as required. Zhang and Lauw (2022) focus on the dynamic topics of temporal document networks and incorporate the linking information between documents. Miyamoto et al. (2023) propose to employ a self-attention mechanism to model the dependencies among dynamic topics. Rather than modeling topic evolution, Cvejoski et al. (2023) model the activities of topics over time. Note that the activities of topics evolve over time but their topics are invariant. Thus, this method does not precisely adhere to the original definition of dynamic topic modeling.
@@ 4.5 Correlated NTMs
Following the idea of correlated topic modeling (Blei and Lafferty, 2006a), correlated NTMs have been explored. Correlated topic models seek to consider the correlation between latent topics. For example, a document about genetics is more likely to be also about disease than x-ray astronomy (Blei and Lafferty, 2006a). This leads to better expressiveness than LDA. Liu et al. (2019) follow the VAE-based NTM and use centralized transformation flow to capture topic correlations. To effectively infer the transformation flow, they present the transformation flow lower bound to regulate the KL divergence term.

### v2

@@ 3.3 NTMs with Metadata
While common NTMs learn topics in an unsupervised manner (only using documents), NTMs can also leverage the metadata of documents to guide topic modeling, similar to supervised LDA Mcauliffe and Blei (2007). In detail, Card et al. (2018) introduce Scholar, a NTM that can incorporate various metadata of documents. It encodes a document together with its labels (e.g., sentiment) and covariates (e.g., publication year), and generates the document conditioned on the covariates. Korshunova et al. (2019) model the generation of documents and labels together in a discriminative way; then train their model with mean-field variational inference. They can also incorporate a variety of data modalities like images. Wang and Yang (2020) jointly model topics and train a RNN classifier to predict document labels. They are connected by an attention mechanism. Wang et al. (2021a) incorporate document networks in a NTM and jointly reconstruct documents and networks.
@@ 3.6 NTMs with Pre-trained Language Models
Researchers frequently combine NTMs with pre-trained language models. Pre-trained language models based on Transformers Vaswani et al. (2017) have been prevalent in NLP fields, which are pre-trained on large-scale corpora to capture contextual linguistic features. Multiple studies leverage contextual features from these pre-trained models to provide richer information than conventional BoW. For instance, Bianchi et al. (2021a) input the concatenation of BoW and the contextual document embeddings from Sentence-BERT Reimers and Gurevych (2019), and then reconstruct BoW as previous work. Hoyle et al. (2020) propose to distill knowledge from BERT Devlin et al. (2018) to NTMs. In detail, they produce pseudo BoW from the predictive word probability of BERT. Then their NTM reconstructs both the real and pseudo BoW. Bianchi et al. (2021b); Mueller and Dredze (2021) employ multilingual BERT to infer cross-lingual doc-topic distributions for zero-shot learning but they cannot discover aligned cross-lingual topics.
@@ 3.7 NTMs with Contrastive Learning
As a self-supervised learning fashion, contrastive learning has been employed to facilitate NTMs Hadsell et al. (2006); Nguyen et al. (2022, 2024a). The idea of contrastive learning is to measure the similarity relations among sample pairs in a representation space Van den Oord et al. (2018). Nguyen and Luu (2021) propose the contrastive learning on doc-topic distributions where they build positive and negative pairs by sampling salient words of documents. Differently, Wu et al. (2022) directly sample positive and negative pairs based on the topic semantics of documents to capture relations among samples. Specifically, they quantize doc-topic distributions following Wu et al. (2020b) and then sample documents with the same quantization indices as positive pairs and different indices as negative pairs. Their method can also capture the similarity relations among additional augmented data. Zhou et al. (2023) improve topic disentanglement with contrastive learning on word and topic embeddings. Han et al. (2023) cluster documents, compute term weights, and make NTMs reconstruct salient words. They also use contrastive learning to refine doc-topic distributions where positive samples come from pre-trained language models. Besides document-level contrastive learning, Nguyen et al. (2024b) also consider topic-level and propose a multi-objective contrastive learning method.
@@ 3.9 Other NTMs
Before the invention of VAE-based NTMs, researchers have different attempts to model latent topics with neural networks. Some studies focus on NTMs in the autoregressive framework. Larochelle and Lauly (2012) propose an autoregressive NTM, called DocNADE. Inspired by Replicated Softmax Hinton and Salakhutdinov (2009), DocNADE predicts the probability of a word in a document conditioned on its hidden state which is conditioned on previous words. Then it interprets topics with a hidden state and infers doc-topic distributions with the hidden states of the document. Gupta et al. (2019b) extend DocNADE by modeling the bi-directional dependencies between words. Gupta et al. (2019a) then use a LSTM to enable DocNADE to incorporate external knowledge.
Cao et al. (2015) also propose an early NTM before VAE-based NTMs. Their approach predicts how an n-gram correlates with documents. It computes the representation of an n-gram by transforming the accumulation of the word embeddings from Word2Vec Mikolov et al. (2013) and projects documents into representations with a look-up matrix table. In this way, they model topic-word distributions as the n-gram representations and model doc-topic distributions as the projected document representations. For training, it uses the document of the n-gram as a positive and randomly samples documents that do not contain this n-gram as negatives.
@@
Pham et al. (2024) propose TopicGPT, prompting Large Language Models (LLMs) to augment seed topics. It defines each topic as a textual description, rather than the word distributions in LDA. It also prompts LLMs to assign a single or more topics to a document, like a classification task. We emphasize that TopicGPT cannot give the topic-word distributions or doc-topic distributions for downstream applications.
@@ 3.10 Topic Discovery by Clustering
We discuss a special type of approach that discovers latent topics by clustering instead of modeling the generation process of documents. They typically leverage traditional word embeddings such as Word2Vec Mikolov et al. (2013) or contextual embeddings from pre-trained language models. We must clarify that some of these methods differ from the aforementioned ordinary topic models. This is because they can only produce topics but cannot infer doc-topic distributions as required. Accordingly, their one advantage is their enhanced computational efficiency. In detail, Thompson and Mimno (2020) straightforwardly cluster token-level word embeddings from pre-trained models like BERT and GPT-2 and produce topics from the words assigned to clusters. Similarly, Sia et al. (2020); Zhang et al. (2022c) cluster word embeddings and interpret hidden topics by sampling words from clusters via term weights like TF-IDF. Angelov (2020) propose Top2Vec. It obtains document and word embeddings through Doc2Vec Le and Mikolov (2014). Then it reduces the dimension of document embeddings with UMAP and clusters them with HDBSCAN. It extracts the n-closest words of a cluster to represent a topic. Following Top2Vec, Grootendorst (2022) propose BERTopic. BERTopic extracts words of a clustering based on c-TF-IDF, which calculated the TF-IDF of a word over a cluster. It estimates the doc-topic distribution based on the term weights within a given document.
3.11 NTM with Semantic-Relations
Wu et al. (2024b) propose FASTopic following a new paradigm rather than VAE-based or clustering-based ones. Similar to Top2Vec and BERTopic, FASTopic leverages document embeddings, like from pretrained Transformers Reimers and Gurevych (2019); Pan et al. (2023); Wu et al. (2024d). Differently, it conducts optimal transport Peyré et al. (2019) between document and topic embeddings and uses the transport plans to model doc-topic distributions. In the same way, it models topic-word distributions with the transport plans between topic and word embeddings. Then it optimizes topic and word embeddings by reconstruction with these semantic relations. This avoids the previous complicated VAE or simple clustering structures, leading to a neat and efficient paradigm.
@@ 4.1 Hierarchical NTMs
Similar to conventional topic models Griffiths et al. (2003); Teh et al. (2004); Blei et al. (2010), NTMs can discover hierarchical topics to reveal topic structures from general to specific. Topics at each level in a hierarchy cover different semantic granularity: child topics tend to be more specific to their parent topics. As shown in Figure 4, a topic about “sports” can derive more specific child topics, like “soccer”, “basketball”, and “tennis”; a topic about “computer” also has specific child topics like “linux”, “programming”, and “windows”. In addition, hierarchical topic modeling can relieve the challenge of determining the number of topics to some extent Blei et al. (2010).
To discover hierarchical topics, some NTMs follow the previous non-parametric setting which allows topic hierarchies to grow dynamically. Isonuma et al. (2020) propose a tree-structured neural topic model with two doubly-recurrent neural networks over the ancestors and siblings respectively Alvarez-Melis and Jaakkola (2017). Note that the tree structure is unbounded, i.e., it can be dynamically updated in a heuristic way during training. Pham and Le (2021) follow this spirit and jointly handle hierarchical topics and document visualization. Chen et al. (2021b) leverage a stick-breaking process as prior for non-parametric modeling.
Later, the parametric fashion has attracted more attention, which presets the width and height of a topic hierarchy before learning. This leads to the main issue that topic hierarchies cannot grow dynamically. Chen et al. (2021a) propose manifold regularization on topic hierarchy learning. Duan et al. (2021) propose a Sawtooth Connection to model topic dependencies across hierarchical levels based on the model structure of ETM Dieng et al. (2020). As aforementioned, Xu et al. (2022) use different layers in the hyperbolic embedding space to interpret hierarchical topics. Li et al. (2022) employ skip-connections for decoding to alleviate the posterior collapsing issue and propose a policy gradient method for training. Recently, Duan et al. (2023) propose to generate different documents for different levels. They craft documents with more related words through word similarity matrices for higher levels, and then progressively generate these documents at each level. Chen et al. (2023) utilize a Gaussian mixture prior and nonlinear structural equations to model topic dependencies between hierarchical levels. Wu et al. (2024c) propose TraCo. It produces sparse and balanced dependencies by modeling them as the transport plan solutions of specially defined optimal transport problems between hierarchical topics. They then introduce a context-aware disentangled decoder to decode documents with topics at each level, which better distributes different semantic granularity to different levels.
@@ 4.2 Short Text NTMs
Researchers apply NTMs to discover topics from short texts. Short texts, prevalent on the Internet in various forms such as tweets, comments, and news headlines, serve as a common medium for individuals to express ideas, comments, and opinions. However, normal topic models often struggle to effectively handle short texts. The principal reason is that topic models depend on the word co-occurrence information to infer latent topics, but such information is extremely sparse in short texts due to their limited context. This challenge, referred to as data sparsity Yan et al. (2013); Wu and Li (2019), hinders topic models from discovering high-quality topics and thus has attracted considerable attention in the research community.
Several studies are proposed to overcome this data sparsity challenge. Lin et al. (2020) use the Archimedean copulas to regularize the discreteness of topic distributions of short texts. Wu et al. (2020b) propose NQTM, which quantizes doc-topic distributions of short texts to quantization vectors following the idea of Van den Oord and Vinyals (2017). By carefully initializing the quantization vectors, it can produce sharper doc-topic distributions that better fit short texts with limited context. They also propose a negative sampling decoder to avoid repetitive topics besides the negative log-likelihood. To address the data sparsity issue, Wang et al. (2021b) use word co-occurrence and semantic correlation graphs to enrich the learning signals of short texts. Zhao et al. (2021d) propose to incorporate entity vector representations into a NTM for short texts. They learn entity vector representations from manually edited knowledge graphs. Based on NQTM Wu et al. (2020b), Wu et al. (2022) further propose TSCTM, a contrastive learning method according to the topic semantics of short texts, which better captures the similarity relations among them. This refines the representations of short texts and thus their doc-topic distributions. They can also adapt to using data augmentation to further mitigate the data sparsity problem.
@@ 4.3 Cross-lingual NTMs
Cross-lingual NTMs are proposed following cross-lingual topic models Mimno et al. (2009). Cross-lingual topic models aim to discover aligned topics in different languages. As exemplified in Figure 5, English and Chinese Topic#3 both refer to “music”, and English and Chinese Topic#5 refer to “celebrity”. In addition if two documents in different languages contain similar latent topics, their inferred doc-topic distributions should be similar. For instance, the doc-topic distributions of the parallel English and Chinese documents in Figure 5 are similar. These aligned cross-lingual topics can reveal commonalities and differences across languages and cultures, which enables cross-lingual text analysis without supervision.
Wu et al. (2020a) propose the first neural cross-lingual topic model, NMTM. It transforms the topic-word distribution to the vocabulary space of another language. Thus the topic-word distributions of one language can incorporate the semantics of another language, which aligns cross-lingual topics. They show that their model outperforms traditional multilingual topic models Shi et al. (2016); Yuan et al. (2018). Later, Wu et al. (2023a) propose InfoCTM. It aligns cross-lingual topics from the perspective of mutual information. This can properly align cross-lingual topics and prevent degenerate topic representations. To address the low-coverage dictionary issue, they also propose a cross-lingual vocabulary linking method that finds more linked words for topic alignment beyond the given dictionary. Bianchi et al. (2021b); Mueller and Dredze (2021) directly learn cross-lingual doc-topic distributions with multilingual BERT. But we emphasize that they cannot discover aligned cross-lingual topics as required.
@@ 4.4 Dynamic NTMs
Dynamic NTMs are explored following dynamic topic models Blei and Lafferty (2006b); Wang et al. (2008). Previous static topic models implicitly assume that documents are exchangeable. However, this assumption is inappropriate since documents are produced sequentially, such as scholarly journals, emails, and news articles. As such, dynamic topic models are proposed. While topics in previous methods are all static, dynamic topic models allow topics to shift over time to capture the topic evolution in sequential documents.
@@
Dieng et al. (2019) first propose a neural dynamic topic model, DETM (Dynamic Embedding Topic Model). It uses word and topic embeddings to interpret latent topics following Dieng et al. (2020) and chains topic embeddings at slice $t$ with topic embeddings at slice $t-1$ by Markov chains. Besides, it uses a LSTM to learn temporal priors of doc-topic distributions. Rahimi et al. (2023) discover topic evolution by clustering documents but cannot infer doc-topic distributions as required. Zhang and Lauw (2022) focus on the dynamic topics of temporal document networks and incorporate the linking information between documents. Following DETM, Miyamoto et al. (2023) propose to employ a self-attention mechanism to model the dependencies among dynamic topics. Wu et al. (2024a) focus on the unassociated topic and repetitive topic issues. Instead of the previous Markov chains fashion, they propose CFDTM with a contrastive learning method to resolve these issues and track topic evolution. Rather than modeling topic evolution, Cvejoski et al. (2023) model the activities of topics over time. Note that the activities of topics evolve over time but their topics are invariant. Thus, this method does not precisely adhere to the original definition of dynamic topic modeling.
@@ 4.5 Correlated NTMs
Following the idea of correlated topic modeling Blei and Lafferty (2006a), correlated NTMs have been explored. Correlated topic models seek to consider the correlation between latent topics. For example, a document about genetics is more likely to be also about disease than x-ray astronomy Blei and Lafferty (2006a). This leads to better expressiveness than LDA. Liu et al. (2019) follow the VAE-based NTM and use centralized transformation flow to capture topic correlations. To effectively infer the transformation flow, they present the transformation flow lower bound to regulate the KL divergence term.